OpenStudy (anonymous):

tan^2 u = (1-cos2u)/(1+cos2u) a) use the identities to prove the power-reducting identity b) explain why this identity does not imply that tan u = square root of (1-cos2u)/(1+cos2u)

OpenStudy (anonymous):

reducing**

OpenStudy (anonymous):

sen^2(x) = 1/2 - 1/2 cos(2x) cos^2(x) = 1/2 + 1/2 cos(2x) so: tan^2 x=1/2 - 1/2 cos(2x)/ 1/2 + 1/2 cos(2x)

OpenStudy (anonymous):

by sen do u mean sin?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

if u know what sin^2(u) and cos^2(u) are, u dviide the two

OpenStudy (anonymous):

but u dont have sin^2 and cos^2

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

tan^2(u) = sin^2(u)/cos^2(u)

OpenStudy (anonymous):

you asking to reduse the power.... so now there is no squares

OpenStudy (anonymous):

after u say sin/cos though what do u do

OpenStudy (anonymous):

oh never mind i get it. how do u answer the second part of question

OpenStudy (anonymous):

tan x= sinx/cosx tan^2x=sin^2x/cos^2x

OpenStudy (anonymous):

thats for the second part??

OpenStudy (anonymous):

no

OpenStudy (anonymous):

then what is part b?

OpenStudy (anonymous):

not sure about that

OpenStudy (anonymous):

oh . sheena u still there? u know the answer to b?

OpenStudy (anonymous):

maybe @experimentX can help, :)

OpenStudy (anonymous):

part b

OpenStudy (experimentx):

not quite sure .. i'll try.

OpenStudy (anonymous):

thanks u have an idea?

OpenStudy (anonymous):

i dont think it has to be that complicated tho dont think too hard

OpenStudy (experimentx):

@myko have any idea??

OpenStudy (anonymous):

maybe becouse of period difference of cos and tan?

OpenStudy (experimentx):

I don't think it should make difference .... i was rather thinking this what if u=0

OpenStudy (anonymous):

ya, that looks more scary....:)

OpenStudy (experimentx):

the right hand side does not equal the left hand side at the multiples of pi. I think this could be the case.

OpenStudy (anonymous):

pi/2

OpenStudy (experimentx):

Oo.. what was i thinking.

OpenStudy (anonymous):

at pi/2 and 3pi/2 ?

OpenStudy (experimentx):

looks like I'm out of my mind ...

OpenStudy (anonymous):

huh

OpenStudy (experimentx):

the left hand side is always positive, while the right hand side in not.

OpenStudy (anonymous):

ya, i think that's it

OpenStudy (experimentx):

it's because we are taking the positive root of a quadratic equation. It should be rather defined as tan x = + (1-cos2u)/(1+cos2u) for 0 < u < pi/2 (in interval that is odd multiple of pi/2) tan x = - (1-cos2u)/(1+cos2u) for pi/2 < u < pi (in interval that is even multiple of pi/2) Well, that solves it.